# Recent Advances of Foundation Language Models-based Continual Learning: A Survey

### v1

@@ 2.2. Continual Learning for Computer Vision
In the realm of computer vision, De et al. (De Lange et al., 2021) address the pressing challenge of continual learning, specifically in the task-incremental setting, where tasks arrive sequentially with clear boundaries. They introduce a novel stability-plasticity trade-off framework tailored for continual learners and undertake a comprehensive experimental analysis, comparing the efficacy of 11 methodologies across three benchmarks. Qu et al. (Qu et al., 2021) present a comprehensive examination of continual learning, highlighting its vital role in the accumulation of knowledge from sequential data streams. This research investigates a range of approaches, encompassing regularization, knowledge distillation, memory-based methods, and more. These approaches are systematically categorized by their characteristics and applications in computer vision. Moreover, Mai et al. (Mai et al., 2022) focus on the realm of online continual learning within image classification, addressing catastrophic forgetting. This study evaluates the efficacy of state-of-the-art methodologies across diverse memory and data configurations. Masana et al. (Masana et al., 2022) present a comprehensive performance evaluation of class-incremental methods applied to image classification. The experiments span a range of scenarios, incorporating large-scale datasets and varied network architectures. Belouadah et al. (Belouadah et al., 2021) pay more attention to class-incremental learning algorithms for visual tasks. This study defines the essential properties of incremental learning algorithms, offers a unified formalization of the class-incremental learning problem, and provides an evaluation framework for detailed analysis.
@@ 2.3. Continual Learning for NLP
Biesialska et al. (Biesialska et al., 2020) address the challenge of continual learning within Natural Language Processing (NLP), wherein conventional architectures struggle to accommodate new tasks without compromising previously acquired knowledge. The research presents an extensive review of methods, including rehearsal, regularization, and architectural approaches, all designed to mitigate the aforementioned challenge. In a similar vein, Ke et al. (Ke and Liu, 2022) offer a focused survey on continual learning within the NLP domain, providing a comprehensive examination of various continual learning settings, methodologies, and challenges. This work presents an in-depth analysis of state-of-the-art approaches and extends original CL settings to be more general and up-to-date. Additionally, it emphasizes the significance of knowledge transfer within NLP and the challenges posed by inter-task class separation.
@@ 2.4. Continual Learning for Other Domains
Recent surveys, surveys like (Zhang and Kim, 2023; Shaheen et al., 2022; Lesort et al., 2020) investigate the advancements in incremental learning for neural recommendation systems and continual learning (CL) in robotics, respectively. Zhang et al. (Zhang and Kim, 2023) make a notable contribution to narrow the gap between academic research and industrial applications through the introduced Incremental Update Recommendation Systems (IURS). They highlight the imperative for real-time updates with streaming data and focus on the distinctive challenges posed by IURS in contrast to traditional Batch Update Recommendation Systems (BURS) and offer a thorough examination of existing literature and evaluation methodologies in this domain. Shaheen et al. (Shaheen et al., 2022) provide a comprehensive overview of contemporary approaches for CL within real-world contexts. Their analysis focuses on learning algorithms that efficiently handle large sequential datasets within computational and memory constraints. The survey also explores challenges in applying CL to autonomous systems, comparing methods across metrics such as computational efficiency, memory utilization, and network complexity. In the field of robotics, it is essential for agents to adapt and interact with their environment using a continuous stream of observations. Thus, Lesort et al. (Lesort et al., 2020) explore CL within this domain, defining CL as a paradigm where both data distribution and learning objectives evolve dynamically. They emphasize the challenges in evaluating CL algorithms in robotic applications and introduce a novel framework alongside metrics tailored to effectively present and assess CL methodologies.
This paper centers on the crucial advancements in CL as applied to foundational language models, which have obtained significant success in the fields of NLP and multimodal. We categorize existing works into offline and online CL based on PLMs, LLMs, and VLMs.
@@ 3.2.1. Offline Continual Learning
Offline CL (Figure 3) comprises three principal scenarios, each distinguished by distinct characteristics: Domain-Incremental Learning, Task-Incremental Learning, and Class-Incremental Learning.
@@
In summary, Domain-Incremental Learning concentrates on adapting the model to the shifts in input data distributions while maintaining consistency in tasks and classes. Task-Incremental Learning necessitates the model’s ability to learn and retain task-specific knowledge over successive tasks. On the other hand, Class-Incremental Learning highlights the gradual integration of new classes into the model’s recognition capabilities without compromising knowledge of previously learned classes.
@@ 3.2.2. Online Continual Learning
In online continual learning (See Figure 4), existing research is categorized into two configurations based on the arrival pattern of tasks: ”Hard Task Boundary” and ”Blurry Task Boundary”:
@@ 4.1.1. PLMs-based DIL
Continual Learning methodologies are frequently employed in the context of Pre-trained Language Models (PLMs), encompassing approaches such as replay-based, regularization-based, and parameter-isolation-based algorithms. Li et al. (Li et al., 2022a) present a regularization-centric framework for Lifelong Learning (LLL) termed RMR-DSE, specifically tailored for sequential operation across multiple domains. Unlike conventional strategies requiring incremental memory allocation, RMR-DSE employs a recall optimization mechanism, driven by regularization, to selectively retain important parameters from preceding tasks. Furthermore, it incorporates a domain drift estimation algorithm to address embedding space shifts. Castellucci et al. (Castellucci et al., 2021) propose a Knowledge Distillation-based Continual Learning (CL-KD) approach with a Teacher-Student framework. When the student model is trained on a new language, the teacher model also transfers its knowledge of supported languages to the student model. Lee et al. (Lee, 2017) introduce a domain-agnostic framework that employs a two-phase training process, initially using synthetic data to learn general conversational patterns and subsequently using human-computer dialogs in customer support. Moreover, the Adaptive Elastic Weight Consolidation algorithm is applied to adjust the loss function, ensuring the balance between acquiring new knowledge and retaining previously learned information.
Gururangan et al. (Gururangan et al., 2022) develop a parameter-isolation-based method, named DEMIX layers, which consists of a collection of experts. The dynamic addition or removal of these experts can enhance the model’s ability to adapt to new domains while maintaining robust performance in previously established ones. PlugLM (Cheng et al., 2023b) is a pre-training model equipped with a differentiable plug-in memory (DPM) designed for domain-adaptive continual training. The core concept behind this approach is to separate the knowledge storage from model parameters using an adaptable key-value memory structure. It enables the explicit retrieval and utilization of knowledge stored within the DPM.
(a) CLASSIC
@@
Figure 5. Frameworks in DIL: CLASSIC (PLM-based) (Ke et al., 2021b), Lifelong-MoE (LLM-based) (Chen et al., 2023), S-Prompts (VLM-based) (Wang et al., 2022b).
@@
Continual domain-adaptive pre-training (DAP-training) (Ke et al., 2023) is based on two main ideas: (1) the general knowledge in the LM and the knowledge gained from prior domains are crucial to mitigating CF and enhancing cross-task knowledge transfer. This is achieved through soft-masking units based on their importance, and (2) the model is designed to develop complementary representations of both the current domain and prior domains, thereby facilitating the integration of knowledge. The key novelty of Continual DAP-training is a soft-masking mechanism that directly controls the update to the LM. Cossu et al. (Cossu et al., 2022) formalize and explore the dynamics of continual pre-training (Pretr) scenarios across language and vision domains. In this framework, models undergo continuous pre-training on a sequential stream of data before subsequent fine-tuning for various downstream tasks.
@@
The adapter architecture incorporates a skip-connection to minimize the number of parameters. A notable exemplar of this approach is AdapterCL (Madotto et al., 2020), which employs residual adapters tailored specifically for task-oriented dialogue systems. This framework, comprising 37 domains, is structured to facilitate continual learning across four important aspects: intent recognition, state tracking, natural language generation, and end-to-end processing. In a related vein, Ke et al. (Ke and Xu, 2021) introduce B-CL model to tackle critical challenges in CL for Aspect-Based Sentiment Classification. B-CL integrates continual learning adapters within capsule network architectures. Aiming at mitigating catastrophic forgetting during fine-tuning, the CLASSIC model (Ke et al., 2021b) presents an innovative solution by deploying adapters to tap into BERT’s capabilities (Figure 5(a)). A novel contrastive continual learning strategy is used to facilitate the transfer of knowledge across tasks and distill insights from previous tasks to subsequent ones. It also effectively eliminates the necessity for task identifiers during testing. Furthermore, Continual PostTraining (CPT) (Ke et al., 2022) introduces two continual learning plug-in modules, termed CL-plugins, embedded within each transformer layer of RoBERTa.
Prompt tuning (Liu et al., 2022), or P-tuning, introduces trainable continuous prompts into the sequence of input word embeddings, while the language model remains frozen. To tackle the challenge of CL under limited labeled data, Qin et al. (Qin and Joty, 2022b) propose a Lifelong Few-shot Language Learning framework (LFPT5). In this framework, prompt tuning, replay and regularization strategies are leveraged. When presented with a new task, the model generates pseudo-labeled samples representative of prior domains. The training process then incorporates these pseudo-labeled samples alongside new task-specific data. Additionally, the KL divergence loss is employed to maintain label consistency between the previous and the current model. Furthermore, Zhu et al. (Zhu et al., 2022) introduced Continual Prompt Tuning (C-PT) as a methodology to address the challenges of continual learning within dialogue systems. C-PT facilitates knowledge transfer between tasks through continual prompt initialization, query fusion, memory replay, and a memory-guided technique.
@@
Instruction tuning-based methods involve transforming a given task into natural language instructions. Qin et al. (Qin et al., 2022) propose ELLE, a novel approach aimed at effectively incorporating continuously expanding streaming data into pre-trained language models (PLMs). It consists of two fundamental components: (1) function-preserved model expansion, which enhances knowledge acquisition efficiency by changing the width and depth of an existing PLM, and (2) pre-trained domain prompts, which significantly enhance the adaptation for downstream tasks by effectively segregating the diverse knowledge acquired during pre-training phases.
@@ 4.1.2. LLMs-based DIL
In many practical scenarios, retraining Language Models (LMs) is challenging due to resource constraints and data privacy concerns. Zhang et al. (Zhang et al., 2024) introduce Continual Proximal Policy Optimization (CPPO) to address this issue. CPPO integrates sample-wise weighting into the Proximal Policy Optimization (PPO) algorithm, effectively balancing policy learning and knowledge retention. Zhang et al. (Zhang et al., 2023b) propose Continual Optimal Policy Regularization (COPR), which calculates the optimal policy distribution without the partition function and uses the previous optimal policy to regularize the current policy. Sun et al. (Sun et al., 2020a) introduce LAMOL, which generates pseudo-samples from previous tasks while training on a new task. It effectively mitigates knowledge loss without requiring additional memory or computational resources. Building on this framework, Wang et al. (Wang et al., 2022a) developed RVAE_LAMOL, which integrates a residual variational autoencoder (RVAE) to encode input data into a unified semantic space, thereby enhancing task representation. This model also incorporates an identity task to enhance the model’s discriminative ability for task identification. To enhance training efficacy, the Alternate Lag Training (ALT) is devised to segment the training process into multiple phases.
To reduce hallucinations in specialized domains such as the Chinese legal domain, Zhang et al. (Zhang et al., 2023c) propose a novel domain adaptation framework, named Adapt-Retrieve-Revise (ARR). It consists of three steps: adapting a 7-billion-parameter language model for initial responses, retrieving corroborative evidence from an external knowledge base, and integrating these to refine the final response with GPT-4. Gogoulou et al. (Gogoulou et al., 2023) study the pros and cons of updating a language model when new data comes from new languages – the case of continual learning under language shift. They feed various languages into the model to examine the impact of pre-training sequence and linguistic characteristics on both forward and backward transfer effects across three distinct model sizes. A new continual learning (CL) problem, named Continual Knowledge Learning (CKL), is introduced by Jang et al. (Jang et al., 2022b). To assess CKL approaches, the authors establish a benchmark and metric measuring knowledge retention, updating, and acquisition.
@@
LLMs have demonstrated remarkable proficiency in tackling open-domain tasks. However, their application in specific domains faces notable challenges, encompassing the lack of domain-specific knowledge, limited capacity to utilize such knowledge, and inadequate adaptation to domain-specific data formats. To address these issues, researchers have explored a novel approach known as continual pre-training, aiming to adapt LLMs to specific domains (Xie et al., 2023; Ma et al., 2023; Cheng et al., 2023a). Among these studies, Cheng et al. (Cheng et al., 2023a) draw inspiration from human learning patterns to develop a novel method that transforms raw corpora into reading comprehension texts. Furthermore, they discover that while training directly on raw data enhances the model’s domain knowledge, it significantly hurts the question-answering capability of the model.
Domain-adaptive Continual Pre-training (DACP) uses a large domain corpus, leading to high costs. To reduce these, Xie et al. (Xie et al., 2023) propose two strategies: Efficient Task-Similar Domain-Adaptive Continual Pre-training (ETS-DACP) and Efficient Task-Agnostic Domain-Adaptive Continual Pre-training (ETA-DACP). ETS-DACP is tailored to improve performance on specific tasks, building task-specific foundational LLMs. Conversely, ETA-DACP selects the most informative samples across the domain. Given the high cost of training LLMs from scratch and limited annotated data in certain domains, Ma et al. (Ma et al., 2023) propose a novel model called EcomGPT-CT. It employs a fusion strategy to exploit semi-structured E-commerce data. Moreover, multiple tasks are designed to assess LLMs’ few-shot in-context learning ability and zero-shot performance after fine-tuning.
@@
Chen et al. (Chen et al., 2023) present an innovative Lifelong Learning framework, termed Lifelong-MoE, which leverages a Mixture-of-Experts (MoE) architecture (Figure 5(b)). This architecture enhances the model’s capacity by incorporating new experts, where previously trained experts and gating mechanisms are frozen.

### v2

@@ 2.2. Continual Learning for Computer Vision
In the realm of computer vision (CV), De et al. (De Lange et al., 2021) concentrate on task incremental classification and offer significant contributions including a taxonomy, a new framework for balancing stability and plasticity, and an extensive comparison of continual learning methods against baselines. Qu et al. (Qu et al., 2021) present a comprehensive examination of continual learning, highlighting its vital role in the accumulation of knowledge from sequential data streams. This research explores multiple methods including regularization, knowledge distillation, memory-based approaches, and more, categorizing them by their characteristics and applications in CV. Moreover, Mai et al. (Mai et al., 2022) focus on the realm of online continual learning within the image classification task. This study evaluates the efficacy of state-of-the-art methods across diverse memory and data configurations. Masana et al. (Masana et al., 2022) conduct a comprehensive evaluation of class-incremental methods for image classification, involving large-scale datasets and various network architectures. Belouadah et al. (Belouadah et al., 2021) focus more on class-incremental learning algorithms specifically for visual tasks, defining key properties of these algorithms, formalizing the class-incremental learning problem, and providing an evaluation framework for thorough analysis.
@@ 2.3. Continual Learning for NLP
Biesialska et al. (Biesialska et al., 2020) address the challenge of continual learning within Natural Language Processing (NLP), wherein conventional architectures struggle to accommodate new tasks without compromising previously acquired knowledge. In a similar vein, Ke et al. (Ke and Liu, 2022) offer a focused survey on continual learning within the NLP domain, providing a comprehensive examination of various continual learning settings, methodologies, and challenges. This work presents an in-depth analysis of state-of-the-art approaches and extends original CL settings to be more general and up-to-date. Gogoulou et al. (Gogoulou et al., 2023) study the pros and cons of updating a language model when new data comes from new languages – the case of CL under language shift. They feed various languages into the model to examine the impact of pre-training sequence and linguistic characteristics on both forward and backward transfer effects across three distinct model sizes.
@@ 2.4. Continual Learning for Other Domains
Recent surveys, surveys like (Zhang and Kim, 2023; Shaheen et al., 2022; Lesort et al., 2020) explore advancements in incremental learning across other domains. Zhang et al. (Zhang and Kim, 2023) focus on neural recommendation systems, particularly introducing the Incremental Update Recommendation Systems (IURS) to bridge the gap between academic research and industrial applications. They discuss the unique challenges of IURS compared to traditional Batch Update Recommendation Systems (BURS) and offer a detailed review of existing literature and evaluation methods. Shaheen et al. (Shaheen et al., 2022) offer a comprehensive overview of the CL methods in real-world contexts, highlighting efficient learning algorithms for handling large sequential datasets within resource constraints and exploring the applicability of CL to autonomous systems. Lesort et al. (Lesort et al., 2020) investigate CL in robotics, defining it as a dynamic learning paradigm where both data distribution and objectives evolve. They also emphasize the challenges in evaluating CL algorithms and introduce a new framework with tailored metrics for assessing CL methods.
Continual learning across domains aims to mitigate catastrophic forgetting by integrating new tasks while preserving knowledge from prior ones (Monaikul et al., 2021; Li et al., 2022b). However, domains like NLP, CV, and robotics exhibit significant differences. CV methods (Wang et al., 2022d, e) typically focus on classification, while NLP involves more complex tasks such as information extraction and text generation. Robotics tasks evolve sequentially, relying on decision-making and state transitions (Kirkpatrick et al., 2017; Aljundi et al., 2018). NLP models, primarily based on PLMs or LLMs, excel in few-shot and transfer learning, whereas CV tasks commonly utilize traditional neural networks (e.g., ResNet). Recently, Vision-Language Models (VLMs) have been introduced in CV to enhance continual learning. Additionally, the pre-trained transformers in NLP enables parameter-efficient strategies like adapter and prompt tuning, allowing adaptation to new tasks with minimal retraining (Ke et al., 2022; Madotto et al., 2020).
This paper centers on the advancements in CL as applied to foundation LMs, which have obtained success in the fields of NLP and multimodal. We categorize existing works into offline and online CL based on PLMs, LLMs, and VLMs.
@@
3. Typical Foundation Language Model
Pre-trained Language Models (PLMs).
PLMs are trained on large-scale textual data using techniques such as Masked Language Modeling (MLM) and Next Sentence Prediction (NSP) for various tasks, including text classification, sentiment analysis, and named entity recognition. Notable examples of PLMs include ELMo (Sarzynska-Wawer et al., 2021), BERT (Devlin, 2018), early versions of the GPT series (Radford et al., 2018), and RoBERTa (Liu et al., 2019a). PLM-based methods aim to mitigate catastrophic forgetting when learning new tasks or domains within a closed-world setting. Continual learning for PLMs focuses on parameter-efficient tuning and replay-based strategies to reduce resource consumption and prevent catastrophic forgetting. Approaches like adapter-based tuning, prompt-based methods, and knowledge distillation are popular because they update the model with few parameters. This ensures that PLMs remain feasible to deploy on standard hardware while keeping computational costs low.
Large Language Models (LLMs).
LLMs are significantly larger in scale than PLMs, often containing billions or even trillions of parameters. These models are trained on vast datasets using language modeling techniques and excel in tasks such as generation, instruction learning, and in-context learning. LLMs are typically exposed to a broader range of complex tasks, from domain-specific corpora to open-domain challenges. Prominent examples include OpenAI’s GPT-3 (Brown, 2020), GPT-4 (Achiam et al., 2023), LLaMA (Touvron et al., 2023), and others. Due to the substantial memory and computational demands associated with LLMs, continual learning strategies for LLMs often focus on optimizing memory efficiency (e.g., orthogonal low-rank adaptation, recyclable tuning) and selectively updating parameters (e.g., Mixture-of-Experts, dynamic task selection), as retraining or fine-tuning the entire model is typically cost-prohibitive. Additionally, LLMs require techniques to mitigate catastrophic forgetting while managing task diversity and retaining foundation knowledge.
Vision-Language Models (VLMs).
VLMs are designed to process and integrate multimodal information, such as text and vision, by building on PLMs or LLMs for tasks like image captioning, visual question answering, and visual reasoning. Notable examples include CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). VLMs must effectively handle multimodal data, which necessitates continual learning methods that can efficiently manage cross-modal learning. Continual learning for VLMs often emphasizes alignment across multiple modalities using specialized modules such as cross-modal fusion techniques, attention mechanisms, and lightweight adapters. Additionally, methods that focus on retaining modality-specific knowledge and enabling parameter-efficient fine-tuning (e.g., Adaptation-CLIP (Liu et al., 2023a), MoE-Adapters4CL (Yu et al., 2024b)) are essential due to the high costs associated with continual learning across both vision and language domains.
@@ 4.2.1. Offline Continual Learning
Offline CL (Figure 3) comprises three principal scenarios, each distinguished by distinct characteristics: Domain-Incremental Learning (e.g., LFPT5 (Qin and Joty, 2022b), LAMOL (Sun et al., 2020a), S-Prompt (Wang et al., 2022b)), Task-Incremental Learning (e.g., PP (Razdaibiedina et al., 2023), Conpet (Song et al., 2023), ZSCL (Zheng et al., 2023)), and Class-Incremental Learning (e.g., EPI (Wang et al., 2023c), IDBR (Huang et al., 2021), GMM (Cao et al., 2024)).
@@
In summary, DIL concentrates on adapting the model to the shifts in input data distributions while maintaining consistency in tasks and classes. TIL necessitates the model’s ability to learn and retain task-specific knowledge over successive tasks. On the other hand, CIL highlights the gradual integration of new classes into the model’s recognition capabilities without compromising knowledge of previously learned classes.
@@ 4.2.2. Online Continual Learning
In online continual learning (Figure 4), the existing researches are categorized into two configurations based on the arrival pattern of tasks: ”Hard Task Boundary” (e.g., MBPA++ (de Masson D’Autume et al., 2019), PEGP (Qiao et al., 2024)) and ”Blurry Task Boundary” (e.g., SIT (Wang et al., 2024a), G-NoCL (Seo et al., 2024)):
@@ 5.1.1. PLMs-based DIL
Continual Learning methods are increasingly used with Pre-trained Language Models (PLMs), adopting various techniques such as replay, regularization, and parameter-isolation. Driven by regularization, RMR_DSE (Li et al., 2022a) optimizes recall and uses domain drift estimation to manage embedding space shifts across domains, effectively reducing catastrophic forgetting in seq2seq tasks under memory constraints. CL-KD (Castellucci et al., 2021) uses a Teacher-Student model to transfer language knowledge efficiently. AEWC (Lee, 2017) is an improved EWC algorithm, where the importance can be computed in an online manner. DEMIX (Gururangan et al., 2022) is a parameter-isolation approach, utilizing a collection of domain-specific experts for dynamic domain adaptation. It offers superior performance in both in-domain and out-of-domain tasks, providing strong generalization to unseen domains and promising improved adaptability and scalability in multi-domain settings. Lastly, PlugLM (Cheng et al., 2023b) (Figure 5(a)) features a differentiable plug-in memory (DPM) for domain-adaptive training, with a flexible key-value structure that decouples knowledge storage from model parameters, enabling more effective knowledge management in domain-adaptive training.
(a) PlugLM
@@
Figure 5. Frameworks in DIL: PlugLM (PLM-based) (Cheng et al., 2023b), Lifelong-MoE (LLM-based) (Chen et al., 2023), S-Prompts (VLM-based) (Wang et al., 2022b).
@@
Pretr (Cossu et al., 2022) formalizes continual pre-training, where models are continuously pre-trained on sequential data streams before fine-tuning for downstream tasks. DAS (Ke et al., 2023) offers an effective solution for continual domain-adaptive pre-training (DAP-training) of LMs. By integrating soft-masking, KL-divergence loss, and contrastive learning, it mitigates catastrophic forgetting and enhances knowledge transfer between domains.
@@
The adapter architecture incorporates a skip-connection to minimize the number of parameters. A notable exemplar of this approach is AdapterCL (Madotto et al., 2020), which employs residual adapters tailored for task-oriented dialogue systems. This makes it ideal for real-world deployment with frequent domain updates without the need for full retraining. In a related vein, B-CL (Ke and Xu, 2021) sets a new standard in Aspect-Based Sentiment Classification by utilizing CL adapters within capsule network architectures, demonstrating that backward knowledge transfer significantly boosts performance while preserving task-specific information. CLASSIC (Ke et al., 2021b) addresses catastrophic forgetting by leveraging adapters in BERT, employing a contrastive CL strategy that facilitates knowledge transfer across tasks without requiring task identifiers during testing. Furthermore, CPT (Ke et al., 2022) incorporates CL-plugins in RoBERTa, effectively improving end-task performance by managing domain-specific knowledge transfer, particularly in scenarios with limited labeled data.
Prompt tuning (Liu et al., 2022), or P-tuning, introduces trainable prompts into the sequence of input word embeddings, while the LM remains frozen. LFPT5 (Qin and Joty, 2022b) integrates prompt tuning, pseudo-labeled samples, and KL divergence to excel in lifelong few-shot language learning (LFLL). The use of pseudo-labeled samples and KL divergence helps maintain label consistency between models, though task transferability requires careful attention. Similarly, C-PT (Zhu et al., 2022) enhances knowledge transfer between tasks in dialogue systems through prompt initialization, query fusion, and memory replay, outperforming baselines like EWC, Replay, and AdapterCL, while being efficient in parameter usage and memory.
@@
Instruction tuning-based methods involve transforming a given task into natural language instructions. ELLE (Qin et al., 2022) incorporates expanding streaming data into PLMs with two key components: (1) function-preserved model expansion, which enhances knowledge acquisition efficiency by changing the width and depth of PLM, and (2) pre-trained domain prompts, which enhance downstream task adaptation by utilizing domain-specific knowledge from the pre-training phase.
@@ 5.1.2. LLMs-based DIL
In Lifelong Language Learning (LLL), a key challenge is training models on sequential NLP tasks while preserving knowledge from previous tasks. LAMOL (Sun et al., 2020a) tackles this by generating pseudo-samples from past tasks during new task training, effectively reducing knowledge loss without additional memory or computational overhead. RVAE_LAMOL (Wang et al., 2022a) builds on this by utilizing a variational autoencoder (RVAE) to map tasks into a unified semantic space, while introducing Alternate Lag Training (ALT) and an identity task to improve stability and task-specific sample generation.
Efficiently and stably aligning with dynamic human preferences is challenging, especially when minimizing retraining costs. CPPO (Zhang et al., 2024b) addresses this issue by integrating sample-wise weighting into the PPO algorithm, effectively balancing policy learning and knowledge retention, particularly in adaptive tasks like summarization and question-answering. Similarly, COPF (Zhang et al., 2023b) reduces reliance on human feedback by regularizing the current policy with past optimal policies. It is well-suited for handling unlabeled data, making it ideal for real-world applications with limited labeled datasets. Furthermore, AMA (Lin et al., 2024) addresses the ”alignment tax”—the trade-off between maximizing alignment rewards and mitigating forgetting—by interpolating between pre- and post-RLHF model weights. This interpolation identifies optimal layer-wise weight combinations, balancing ”alignment tax” across various RLHF algorithms.
@@
Large language models (LLMs) excel in open-domain tasks but face significant challenges in domain-specific applications, such as insufficient domain knowledge, limited capacity to leverage that knowledge, and inadequate adaptation to specialized data formats. To overcome these limitations, continual pre-training (Xie et al., 2023; Ma et al., 2023; Cheng et al., 2023a) has emerged as a promising approach, enabling LLMs to better fit domain-specific tasks. For example, AdaptLLM (Cheng et al., 2023a) transforms raw corpora into reading comprehension texts, but it was found that while pre-training on raw data enhances domain knowledge, it hampers the model’s ability to answer questions effectively. As an alternative to full retraining, Domain-Adaptive Continual Pre-training (DACP) leverages large domain-specific corpora to continually train models, although it involves high computational costs. DACP (Xie et al., 2023) introduces Efficient Task-Specific (ETS-DACP) and Task-Agnostic (ETA-DACP) strategies to optimize performance by focusing either on task-specific foundational models or by selecting the most informative domain samples. Given the high cost of training from scratch and limited annotated data, EcomGPT-CT (Ma et al., 2023) explores continual pre-training on unlabeled general and E-commerce corpora, employing a data-mixing strategy to better handle semi-structured data. Experimental results across multiple tasks show continual pre-training and data-mixing strategies improve few-shot and zero-shot performance.
@@
Lifelong-MoE (Chen et al., 2023) (Figure 5(b)) uses a Mixture-of-Experts (MoE) architecture, expanding capacity by adding new experts while freezing prior experts and gating mechanisms. It surpasses existing methods with better few-shot performance on 19 NLP tasks.
Instruction Tuning-based Methods.
Adapt-Retrieve-Revise (ARR)(Zhang et al., 2023c) aims at reducing hallucinations in specialized domains such as the Chinese legal domain. It consists of three steps: adapting a 7-billion-parameter language model for initial responses, retrieving corroborative evidence from an external knowledge base, and integrating these to refine the final response with GPT-4. Furthermore, LLM-CL (Ding et al., 2024), designed for aspect-based sentiment analysis, employs task-specific prompts and independently learns both domain-shared and domain-specific knowledge.